# An effective Bombieri–Vinogradov error term for sifting problems

### v1

@@ 2.2  Bounds relating to $\psi(x;d,a)$
with $\psi(x)=\psi(x,1,0)$.
We begin with an effective result of Sedunova, which is an averaged bound for the error (2.6) over moduli $Q_{1}\leq d\leq D$. As with Akbary and Hambrook’s result (Theorem 1.4), this result is effective as it avoids the complications due to the Siegel zero at small moduli.
Lemma 2.4 ([33, Corollary 1.4]).
One has,
$\sum_{Q_{1}\leq d\leq D}\sup_{2\leq y\leq\operatorname{\mathcal{X}}}\max_{\begin{subarray}{c}(a,d)=1\end{subarray}}\left|\psi(y;d,a)-\frac{\psi(y)}{\varphi(d)}\right|\ll_{B}\frac{\operatorname{\mathcal{X}}}{(\log\operatorname{\mathcal{X}})^{B-2}},$ (2.7)
@@
Remark.
The statement of [33, Corollary 1.4] is actually weaker than Lemma 2.4. However, (2.7) is what is shown in the proof of [33, Corollary 1.4].
@@
Lemma 2.5 (See [2, Lemma 6.10] or [3, Theorem 1.2]).
Let $d\leq Q_{1}$ and $\beta_{d}$ denote the Siegel zero modulo $d$ if it exists. Then for any $y_{0}\leq y\leq\mathcal{X}$ and $a\in\mathbb{Z}$ with $(a,d)=1$,
$\varphi(d)\left|\psi(y;d,a)-\frac{\psi(y)}{\varphi(d)}\right|\ll_{B}\begin{cases}\frac{y}{(\log y)^{B}}+y^{\beta_{d}},&\text{if}\ \beta_{d}\ \text{exists},\\ \frac{y}{(\log y)^{B}},&\text{otherwise.}\end{cases}$ (2.8)
@@
Throughout we will also make regular use of the following two lemmas. The first is the the Brun–Titchmarsh theorem, given in its standard form for both $\pi$ and $\psi$. The second is a bound for the partial sums of $\mu^{2}(d)\gamma^{\omega(d)}/\varphi(d)$, which unsurprisingly appears when bounding a sieve remainder term of the form (1.14).
Lemma 2.6 (Brun–Titchmarsh theorem).
For any $d<x$ and $a\in\mathbb{Z}$ with $(a,d)=1$, we have
$\pi(x;d,a)\ll\frac{x}{\varphi(d)\log(x/d)}\quad\text{and}\quad\psi(x;d,a)\ll\frac{x\log x}{\varphi(d)\log(x/d)}+\sqrt{x}\log x.$
with each of the implied constants effective.
Proof.
See for example [23, Theorem 2] for the proof of the $\pi(x;d,a)$ bound. For the $\psi(x;d,a)$ bound, we note that
$\psi(x;d,a)\leq\log x\left(\pi(x;d,a)+\sum_{\begin{subarray}{c}p^{m}\leq x\\ m\geq 2\end{subarray}}1\right).$ (2.9)
For the sum over $p^{m}\leq x$ in (2.9), one only needs to consider $m\leq\log x/\log 2$ since otherwise $p^{m}>x$. Hence, by the prime number theorem
$\sum_{\begin{subarray}{c}p^{m}\leq x\\ m\geq 2\end{subarray}}1\leq\frac{\log x}{\log 2}\sum_{p^{2}\leq x}1\sim\frac{2\sqrt{x}}{\log 2}$
and the result follows. ∎
@@
$\sum_{d\leq D}\frac{\mu^{2}(d)\gamma^{\omega(d)}}{\varphi(d)}\leq\prod_{p\leq D}\left(1+\frac{\gamma}{p-1}\right)=\prod_{p\leq D}\left(1+\frac{\gamma}{p}\right)\prod_{p\leq D}\left(1+\frac{\gamma}{(p+\gamma)(p-1)}\right).$
@@
$\prod_{p\leq D}\left(1+\frac{\gamma}{(p+\gamma)(p-1)}\right)=O_{\gamma}(1)$
@@
$\prod_{p\leq D}\left(1+\frac{\gamma}{p}\right)=\exp\left(\sum_{p\leq D}\log\left(1+\frac{\gamma}{p}\right)\right)$ $\ll\exp\left(\sum_{p\leq D}\frac{\gamma}{p}+O_{\gamma}(1)\right)$
$\ll_{h}(\log x)^{\gamma},$ (2.10)
where in the last step we used Mertens’ first theorem. ∎
Equipped with Lemma 2.5 and 2.6, we now prove a complete equidistribution result in the case where the exceptional modulus is small. Namely, if the exceptional zero is small enough then we can readily apply Lemma 2.2 to bound the contribution of the Siegel zero in (2.8).
@@
$\sum_{\begin{subarray}{c}1\leq d\leq D\\ (d,\overline{\operatorname{\mathcal{P}}})=1\end{subarray}}\mu^{2}(d)\sup_{y\leq\operatorname{\mathcal{X}}}\max_{(a,d)=1}\left|\psi(y;d,a)-\frac{\psi(y)}{\varphi(d)}\right|\ll_{B}\frac{\operatorname{\mathcal{X}}}{(\log\operatorname{\mathcal{X}})^{B-2}}$ (2.11)
@@
We only need to consider the sum over $1\leq d\leq Q_{1}$ since the range $Q_{1}\leq d\leq D$ is covered by Lemma 2.4. We consider three cases for $y$.
Case 1: $2\leq y\leq\sqrt{\operatorname{\mathcal{X}}}$.
In this case, we use the trivial bounds $\psi(y;d,a)\leq\psi(y)\ll y\log y$ to obtain
$\sum_{\begin{subarray}{c}1\leq d\leq Q_{1}\\ (d,\overline{\operatorname{\mathcal{P}}})=1\end{subarray}}\mu^{2}(d)\max_{(a,d)=1}\left|\psi(y;d,a)-\frac{\psi(y)}{\varphi(d)}\right|\ll(y\log y)Q_{1}\ll\sqrt{\operatorname{\mathcal{X}}}(\log\operatorname{\mathcal{X}})^{B+1},$
which is sufficient.
Case 2: $\sqrt{\operatorname{\mathcal{X}}}\leq y\leq y_{0}$.
In this case, the Brun–Titchmarsh theorem (Lemma 2.6) gives
$\psi(y;d,a)\ll_{B}\frac{y_{0}}{\varphi(d)}\quad\text{and}\quad\frac{\psi(y)}{\varphi(d)}\ll\frac{y_{0}}{\varphi(d)}.$ (2.12)
for any $1\leq d\leq Q_{1}$. Using (2.12) along with Lemma 2.7 (with $\gamma=1$) then yields
$\sum_{\begin{subarray}{c}1\leq d\leq Q_{1}\\ (d,\overline{\operatorname{\mathcal{P}}})=1\end{subarray}}\mu^{2}(d)\max_{(a,d)=1}\left|\psi(y;d,a)-\frac{\psi(y)}{\varphi(d)}\right|$ $\ll_{B}y_{0}\sum_{d\leq Q}\frac{\mu^{2}(d)}{\varphi(d)}\ll\frac{\operatorname{\mathcal{X}}\log\log\operatorname{\mathcal{X}}}{(\log\operatorname{\mathcal{X}})^{B}},$
which is sufficient.
Case 3: $y_{0}\leq y\leq\operatorname{\mathcal{X}}$.
For this range of $y$, if there does not exist a Siegel zero mod $d$ for all $d\leq Q_{1}$, then the desired result follows by applying Lemma 2.5 pointwise for $d\leq Q_{1}$, followed by Lemma 2.7. On the other hand, if there exists a Siegel zero $\beta_{d}$ induced by a character $\chi_{d}$ mod $d$, then there are three subcases to consider:
@@
Hence, by Lemmas 2.5 and 2.7, we have
$\sum_{\begin{subarray}{c}1\leq d\leq Q_{1}\\ (d,\overline{\operatorname{\mathcal{P}}})=1\end{subarray}}\mu^{2}(d)\max_{(a,d)=1}\left|\psi(y;d,a)-\frac{\psi(y)}{\varphi(d)}\right|$ $\ll_{B}\sum_{\begin{subarray}{c}1\leq d\leq Q_{1}\\ (d,\overline{\mathcal{P}})=1\end{subarray}}\frac{\mu^{2}(d)}{\varphi(d)}\left(\frac{\operatorname{\mathcal{X}}}{(\log\operatorname{\mathcal{X}})^{B}}+\operatorname{\mathcal{X}}^{1-1/\sqrt{\log\operatorname{\mathcal{X}}}}\right)$
@@
$\sum_{\begin{subarray}{c}1\leq d\leq D\\ (d,\overline{\operatorname{\mathcal{P}}})=1\\ k_{1}\nmid d\end{subarray}}\mu^{2}(d)\sup_{y\leq\operatorname{\mathcal{X}}}\max_{(a,d)=1}\left|\psi(y;d,a)-\frac{\psi(y)}{\varphi(d)}\right|\ll_{B}\frac{\operatorname{\mathcal{X}}}{(\log\operatorname{\mathcal{X}})^{B-2}}$
@@
Identical to the proof Proposition 2.8 except we do not include the case $k_{1}\mid d$, which required $k_{1}\leq\log\operatorname{\mathcal{X}}$. ∎
@@ 2.3  Bounds relating to $\pi(x;d,a)$
We now convert our results on $\psi(x;d,a)$ to results in terms of the usual prime-counting function $\pi(x;d,a)$. The proof of the following result is similar to that of [32, Corollary 1.5] but we take more care so as to save an extra factor of $\log\operatorname{\mathcal{X}}$. Moreover, to simplify our argument, we restrict to $y\in[y_{0},\operatorname{\mathcal{X}}]$ for now, but later weaken this restriction in our subsequent, more general result (Proposition 2.12).
@@
$\sum_{\begin{subarray}{c}1\leq d\leq D\\ (d,\overline{\operatorname{\mathcal{P}}})=1\\ k_{1}\nmid d\end{subarray}}\mu^{2}(d)\sup_{y_{0}\leq y\leq\operatorname{\mathcal{X}}}\max_{(a,d)=1}\left|\pi(y;d,a)-\frac{\pi(y)}{\varphi(d)}\right|\ll_{B}\frac{\operatorname{\mathcal{X}}}{(\log\operatorname{\mathcal{X}})^{B-1}}$ (2.13)
@@
$\sum_{\begin{subarray}{c}1\leq d\leq D\\ (d,\overline{\operatorname{\mathcal{P}}})=1\end{subarray}}\mu^{2}(d)\sup_{y_{0}\leq y\leq\operatorname{\mathcal{X}}}\max_{(a,d)=1}\left|\pi(y;d,a)-\frac{\pi(y)}{\varphi(d)}\right|\ll_{B}\frac{\operatorname{\mathcal{X}}}{(\log\operatorname{\mathcal{X}})^{B-1}}$ (2.14)
@@
$\pi_{1}(y):=\sum_{2\leq n\leq y}\frac{\Lambda(n)}{\log n}\quad\text{and}\quad\pi_{1}(y;d,a):=\sum_{\begin{subarray}{c}2\leq n\leq y\\ n\equiv a\text{(mod}\ d)\end{subarray}}\frac{\Lambda(n)}{\log n}.$
@@
$\sum_{\begin{subarray}{c}1\leq d\leq D\\ (d,\overline{\operatorname{\mathcal{P}}})=1\end{subarray}}\mu^{2}(d)\max_{(a,d)=1}\left|\pi_{1}(y;d,a)-\frac{\pi_{1}(y)}{\varphi(d)}\right|\ll_{B}\frac{\operatorname{\mathcal{X}}}{(\log\operatorname{\mathcal{X}})^{B-1}}$ (2.16)
uniformly for $y\in[y_{0},\operatorname{\mathcal{X}}]$. We split the integral in (2.15) into the range $2\leq t\leq y^{0.1}$ and $y^{0.1}\leq t\leq y$. In the first case, the trivial bounds $\psi(t;d,a)\leq\psi(t)\leq t\log t$} give
$\int_{2}^{y^{0.1}}\frac{\psi(t;d,a)-\psi(t)/\varphi(d)}{t(\log t)^{2}}\mathrm{d}t\ll y^{0.1}\int_{2}^{y^{0.1}}\frac{1}{\log t}\mathrm{d}t\ll\operatorname{\mathcal{X}}^{0.2}.$ (2.17)
@@
$\quad\ll\frac{1}{\log y}\left|\psi(y;d,a)-\frac{\psi(y)}{\varphi(d)}\right|+\sup_{\sqrt{y}\leq t\leq y}\left|\psi(t;d,a)-\frac{\psi(t)}{\varphi(d)}\right|\int_{\sqrt{y}}^{y}\frac{1}{t(\log t)^{2}}\mathrm{d}t+\operatorname{\mathcal{X}}^{0.2}$
$\quad\ll\frac{1}{\log\operatorname{\mathcal{X}}}\sup_{\sqrt{y}\leq t\leq y}\left|\psi(t;d,a)-\frac{\psi(t)}{\varphi(d)}\right|+\operatorname{\mathcal{X}}^{0.2}.$
@@
$\pi_{1}(y;d,a)-\pi(y;d,a)=\sum_{2\leq k\leq\frac{\log k}{\log 2}}\sum_{\begin{subarray}{c}p^{k}\leq y\\ p^{k}\equiv a\text{(mod}\ d)\end{subarray}}\frac{1}{k}\leq\sum_{2\leq k\leq\frac{\log k}{\log 2}}\frac{\pi(\sqrt{y})}{2}\ll\sqrt{y},$ (2.18)
where the last inequality follows from the prime number theorem (or the Brun–Titchmarsh theorem). Similarly,
@@
To finish, we substitute (2.18) and (2.19) into (2.16), giving
$\sum_{\begin{subarray}{c}1\leq d\leq D\\ (d,\overline{\operatorname{\mathcal{P}}})=1\end{subarray}}\mu^{2}(d)\max_{(a,d)=1}\left|\pi(y;d,a)-\frac{\pi(y)}{\varphi(d)}\right|\ll_{B}\frac{\operatorname{\mathcal{X}}}{(\log\operatorname{\mathcal{X}})^{B-1}}+\sqrt{y}D\ll\frac{\operatorname{\mathcal{X}}}{(\log\operatorname{\mathcal{X}})^{B-1}}$
@@
We only prove (2.21) as the proof of (2.20) is essentially identical. The case $\gamma\leq 1$ follows immediately from Proposition 2.8 so we assume $\gamma>1$ throughout. As with our previous results, we split into three cases for $y$.
@@
$B-2\gamma+1\geq B_{\gamma}:=(B-\gamma^{2})/2$
for $B>\gamma^{2}$.
@@
In this case, the Brun–Titchmarsh theorem (Lemma 2.6) gives
@@
for all $d\leq D$. Using these bounds and Lemma 2.7, we have
@@
$\sum_{\begin{subarray}{c}d\leq D\\ (d,\overline{\operatorname{\mathcal{P}}})=1\end{subarray}}E(d)\ll_{B}\frac{\operatorname{\mathcal{X}}}{(\log\operatorname{\mathcal{X}})^{B-1}}.$ (2.25)
Next, an application of the Brun–Titchmarsh theorem and Lemma 2.7 gives
@@
As one might expect, a small improvement is possible if one uses Hölder’s inequality in place of the Cauchy–Schwarz inequality. Namely, provided $B>\gamma+2$, one can apply Hölder’s inequality in the proof above, writing
$\sum_{\begin{subarray}{c}d\leq D\\ (d,\overline{\operatorname{\mathcal{P}}})=1\end{subarray}}f(d)E(d)\leq\Bigg({\sum_{\begin{subarray}{c}d\leq D\\ (d,\overline{\operatorname{\mathcal{P}}})=1\end{subarray}}f(d)^{p}E(d)}\Bigg)^{1/p}\Bigg(\sum_{\begin{subarray}{c}d\leq D\\ (d,\overline{\operatorname{\mathcal{P}}})=1\end{subarray}}E(d)\Bigg)^{1/q}.$
@@
$(p,q)=\left(\frac{\log(B-2)}{\log\gamma},\frac{\log(B-2)}{\log(B-2)-\log\gamma}\right).$
@@
$\widetilde{B}_{\gamma}=B-1-\frac{2(B-2)\log\gamma}{\log(B-2)}.$
So, for example, if $B=10$ and $\gamma=2$ then $\widetilde{B}_{\gamma}=11/3$ which is greater than $B_{\gamma}=3$. Nevertheless, we have used $B_{\gamma}$ in the statement of Theorems 1.6 and 1.7 for simplicity.

### v2

@@ 2.2  Bounds relating to $\psi(x;d,a)$
with $\psi(x)=\psi(x;1,0)$.
We begin with an averaged bound for the error (2.7) over moduli $Q_{1}<d\leq D$, which corrects a result of Sedunova [35]. As with Akbary and Hambrook’s result (Theorem 1.4), this result is effective as it avoids the complications due to the Siegel zero at small moduli.
Lemma 2.4 ([35, Corollary 1.4]).
One has
$\sum_{Q_{1}<d\leq D}\sup_{y\leq\operatorname{\mathcal{X}}}\max_{\begin{subarray}{c}(a,d)=1\end{subarray}}\left|\psi(y;d,a)-\frac{\psi(y)}{\varphi(d)}\right|\ll_{B}\frac{\operatorname{\mathcal{X}}}{(\log\operatorname{\mathcal{X}})^{B-3}},$ (2.8)
@@
Proof.
The proof of the lemma is essentially the same as that of [35, Corollary 1.4]. However, we outline how to correct a small error in this result, which ultimately gives the denominator $(\log\operatorname{\mathcal{X}})^{B-3}$ in (2.8) as opposed to $(\log\operatorname{\mathcal{X}})^{B-2}$.
The key tool required is Vaughan’s inequality, in which Sedunova [35, Proposition 1.3] is able to prove in the form
$\sum_{q\leq Q}\frac{q}{\varphi(q)}{\sum_{\chi(q)}}^{*}\sup_{y\leq x}|\psi(y,\chi)|\ll_{\varepsilon}(x+Q^{2}x^{1/2}+Qx^{13/14+\varepsilon})(\log x)^{2}$ (2.9)
for any $Q,\varepsilon>0$. Here, for a Dirichlet character $\chi$
$\psi(y,\chi):=\sum_{n\leq y}\Lambda(n)\chi(y)$
and the starred sum in (2.9) indicates that one is only summing over primitive characters $\chi$ mod $q$. Arguing as in [35, §3], one then has
$\sum_{Q_{1}<d\leq D}\sup_{y\leq\operatorname{\mathcal{X}}}\max_{\begin{subarray}{c}(a,d)=1\end{subarray}}\left|\psi(y;d,a)-\frac{\psi(y)}{\varphi(d)}\right|$
$\qquad\qquad\qquad\qquad\ll\log\operatorname{\mathcal{X}}\sum_{Q_{1}<d\leq D}\frac{1}{\varphi(d)}{\sum_{\chi(q)}}^{*}\sup_{2\leq y\leq\operatorname{\mathcal{X}}}|\psi(y,\chi)|+D(\log\operatorname{\mathcal{X}})^{3}.$
Next, one uses partial summation to get that
$\sum_{Q_{1}<d\leq D}\frac{1}{\varphi(d)}{\sum_{\chi(q)}}^{*}\sup_{2\leq y\leq\operatorname{\mathcal{X}}}|\psi(y,\chi)|=\frac{1}{D}\sum_{q\leq D}h(q)-\frac{1}{Q_{1}}\sum_{q\leq Q_{1}}h(q)+\int_{Q_{1}}^{D}\left(\sum_{q\leq t}h(q)\right)\frac{\mathrm{d}t}{t^{2}},$
where
$h(q):=\frac{q}{\varphi(q)}{\sum_{\chi(q)}}^{*}\sup_{y\leq\operatorname{\mathcal{X}}}|\psi(y,\chi)|.$
Applying (2.9) thus gives
$\sum_{Q_{1}<d\leq D}\sup_{y\leq\operatorname{\mathcal{X}}}\max_{\begin{subarray}{c}(a,d)=1\end{subarray}}\left|\psi(y;d,a)-\frac{\psi(y)}{\varphi(d)}\right|\ll_{\varepsilon}\left(\frac{\operatorname{\mathcal{X}}}{Q_{1}}+D\operatorname{\mathcal{X}}^{1/2}+\operatorname{\mathcal{X}}^{13/14+\varepsilon}\log\left(\frac{D}{Q_{1}}\right)\right)(\log\operatorname{\mathcal{X}})^{3},$
from which (2.8) follows. ∎
@@
Lemma 2.5 (See e.g. [2, Lemma 6.10] or [4, Theorem 1.2]).
Let $d\leq Q_{1}$ and $\beta_{d}$ denote the Siegel zero modulo $d$ if it exists. Then,
$\sup_{y\leq\operatorname{\mathcal{X}}}\max_{(a,d)=1}\left|\psi(y;d,a)-\frac{\psi(y)}{\varphi(d)}\right|\ll_{B}\frac{1}{\varphi(d)}\cdot\begin{cases}\frac{\operatorname{\mathcal{X}}}{(\log\operatorname{\mathcal{X}})^{B}}+\operatorname{\mathcal{X}}^{\beta_{d}},&\text{if}\ \beta_{d}\ \text{exists},\\ \frac{\operatorname{\mathcal{X}}}{(\log\operatorname{\mathcal{X}})^{B}},&\text{otherwise.}\end{cases}$ (2.10)
@@
Throughout we will also make regular use of the following bound for the partial sums of $\mu^{2}(d)\gamma^{\omega(d)}/\varphi(d)$, which unsurprisingly appears when bounding a sieve remainder term of the form (1.15).
@@
$\sum_{d\leq x}\frac{\mu^{2}(d)\gamma^{\omega(d)}}{\varphi(d)}\leq\prod_{p\leq x}\left(1+\frac{\gamma}{p-1}\right)=\prod_{p\leq x}\left(1+\frac{\gamma}{p}\right)\prod_{p\leq x}\left(1+\frac{\gamma}{(p+\gamma)(p-1)}\right).$
@@
$\prod_{p\leq x}\left(1+\frac{\gamma}{(p+\gamma)(p-1)}\right)=O_{\gamma}(1)$
@@
$\prod_{p\leq x}\left(1+\frac{\gamma}{p}\right)=\exp\left(\sum_{p\leq x}\log\left(1+\frac{\gamma}{p}\right)\right)$ $\ll\exp\left(\sum_{p\leq x}\frac{\gamma}{p}+O_{\gamma}(1)\right)$
$\ll_{\gamma}(\log x)^{\gamma},$ (2.11)
where in the last step we used Mertens’ theorem. ∎
Equipped with these preliminary lemmas, we now prove a complete equidistribution result in the case where the exceptional modulus is small. Notably, if the exceptional zero is small enough then we can readily apply Lemma 2.3 to bound the contribution of the Siegel zero in (2.10).
@@
$\sum_{\begin{subarray}{c}1\leq d\leq D\\ (d,\overline{\operatorname{\mathcal{P}}})=1\end{subarray}}\mu^{2}(d)\sup_{y\leq\operatorname{\mathcal{X}}}\max_{(a,d)=1}\left|\psi(y;d,a)-\frac{\psi(y)}{\varphi(d)}\right|\ll_{B}\frac{\operatorname{\mathcal{X}}}{(\log\operatorname{\mathcal{X}})^{B-3}}$ (2.12)
@@
We only need to consider the sum over $1\leq d\leq Q_{1}$ since the range $Q_{1}<d\leq D$} is covered by Lemma 2.4.
First, suppose that there does not exist a Siegel zero mod $d$ for all $d\leq Q_{1}$, then the desired result follows by applying Lemma 2.5 pointwise for $d\leq Q_{1}$, followed by Lemma 2.6. On the other hand, if there exists a Siegel zero $\beta_{d}$ induced by a character $\chi_{d}$ mod $d$, then there are three subcases to consider:
@@
Hence, by Lemmas 2.5 and 2.6, we have
$\sum_{\begin{subarray}{c}1\leq d\leq Q_{1}\\ (d,\overline{\operatorname{\mathcal{P}}})=1\end{subarray}}\mu^{2}(d)\sup_{y\leq\operatorname{\mathcal{X}}}\max_{(a,d)=1}\left|\psi(y;d,a)-\frac{\psi(y)}{\varphi(d)}\right|$ $\ll_{B}\sum_{\begin{subarray}{c}1\leq d\leq Q_{1}\\ (d,\overline{\mathcal{P}})=1\end{subarray}}\frac{\mu^{2}(d)}{\varphi(d)}\left(\frac{\operatorname{\mathcal{X}}}{(\log\operatorname{\mathcal{X}})^{B}}+\operatorname{\mathcal{X}}^{1-1/\sqrt{\log\operatorname{\mathcal{X}}}}\right)$
@@
$\sum_{\begin{subarray}{c}1\leq d\leq D\\ (d,\overline{\operatorname{\mathcal{P}}})=1\\ k_{1}\nmid d\end{subarray}}\mu^{2}(d)\sup_{y\leq\operatorname{\mathcal{X}}}\max_{(a,d)=1}\left|\psi(y;d,a)-\frac{\psi(y)}{\varphi(d)}\right|\ll_{B}\frac{\operatorname{\mathcal{X}}}{(\log\operatorname{\mathcal{X}})^{B-3}}$
@@
Identical to the proof of Proposition 2.7 except we do not include the case $k_{1}\mid d$ (i.e. Case (c)), which required $k_{1}\leq\log\operatorname{\mathcal{X}}$. ∎
@@ 2.3  Bounds relating to $\pi(x;d,a)$
We now convert our results on $\psi(x;d,a)$ to results in terms of the usual prime-counting function $\pi(x;d,a)$. The proof of the following result is similar to that of [34, Corollary 1.5] but we take more care so as to save an extra factor of $\log\operatorname{\mathcal{X}}$. Moreover, to simplify our argument, we restrict to $y\in[y_{0},\operatorname{\mathcal{X}}]$ for now, but later weaken this restriction in our subsequent, more general result (Proposition 2.11).
@@
$\sum_{\begin{subarray}{c}1\leq d\leq D\\ (d,\overline{\operatorname{\mathcal{P}}})=1\\ k_{1}\nmid d\end{subarray}}\mu^{2}(d)\sup_{y_{0}\leq y\leq\operatorname{\mathcal{X}}}\max_{(a,d)=1}\left|\pi(y;d,a)-\frac{\pi(y)}{\varphi(d)}\right|\ll_{B}\frac{\operatorname{\mathcal{X}}}{(\log\operatorname{\mathcal{X}})^{B-2}}$ (2.13)
@@
$\sum_{\begin{subarray}{c}1\leq d\leq D\\ (d,\overline{\operatorname{\mathcal{P}}})=1\end{subarray}}\mu^{2}(d)\sup_{y_{0}\leq y\leq\operatorname{\mathcal{X}}}\max_{(a,d)=1}\left|\pi(y;d,a)-\frac{\pi(y)}{\varphi(d)}\right|\ll_{B}\frac{\operatorname{\mathcal{X}}}{(\log\operatorname{\mathcal{X}})^{B-2}}$ (2.14)
@@
$\pi_{1}(y):=\sum_{2\leq n\leq y}\frac{\Lambda(n)}{\log n}\quad\text{and}\quad\pi_{1}(y;d,a):=\sum_{\begin{subarray}{c}2\leq n\leq y\\ n\equiv a\thinspace\text{(mod}\ d)\end{subarray}}\frac{\Lambda(n)}{\log n}.$
@@
$\sum_{\begin{subarray}{c}1\leq d\leq D\\ (d,\overline{\operatorname{\mathcal{P}}})=1\end{subarray}}\mu^{2}(d)\max_{(a,d)=1}\left|\pi_{1}(y;d,a)-\frac{\pi_{1}(y)}{\varphi(d)}\right|\ll_{B}\frac{\operatorname{\mathcal{X}}}{(\log\operatorname{\mathcal{X}})^{B-2}}$ (2.16)
uniformly for $y\in[y_{0},\operatorname{\mathcal{X}}]$. We split the integral in (2.15) into the ranges $2\leq t\leq y^{0.1}$ and $y^{0.1}\leq t\leq y$. In the first case, the trivial bounds $\psi(t;d,a)\leq\psi(t)\leq t\log t$} give
$\int_{2}^{y^{0.1}}\frac{\psi(t;d,a)-\psi(t)/\varphi(d)}{t(\log t)^{2}}\mathrm{d}t\ll\int_{2}^{y^{0.1}}\frac{1}{\log t}\mathrm{d}t\ll\operatorname{\mathcal{X}}^{0.1}.$ (2.17)
@@
$\quad\ll\frac{1}{\log y}\left|\psi(y;d,a)-\frac{\psi(y)}{\varphi(d)}\right|+\sup_{y^{0.1}\leq t\leq y}\left|\psi(t;d,a)-\frac{\psi(t)}{\varphi(d)}\right|\int_{y^{0.1}}^{y}\frac{1}{t(\log t)^{2}}\mathrm{d}t+\operatorname{\mathcal{X}}^{0.1}$
$\quad\ll\frac{1}{\log\operatorname{\mathcal{X}}}\sup_{y^{0.1}\leq t\leq y}\left|\psi(t;d,a)-\frac{\psi(t)}{\varphi(d)}\right|+\operatorname{\mathcal{X}}^{0.1}.$
@@
$\pi_{1}(y;d,a)-\pi(y;d,a)=\sum_{2\leq k\leq\frac{\log y}{\log 2}}\sum_{\begin{subarray}{c}p^{k}\leq y\\ p^{k}\equiv a\thinspace\text{(mod}\ d)\end{subarray}}\frac{1}{k}\leq\sum_{2\leq k\leq\frac{\log y}{\log 2}}\frac{\pi(\sqrt{y})}{2}\ll\sqrt{y},$ (2.18)
where the last inequality follows from the prime number theorem. Similarly,
@@
To finish, we substitute (2.16), (2.18) and (2.19) into (2.13), giving
$\sum_{\begin{subarray}{c}1\leq d\leq D\\ (d,\overline{\operatorname{\mathcal{P}}})=1\end{subarray}}\mu^{2}(d)\max_{(a,d)=1}\left|\pi(y;d,a)-\frac{\pi(y)}{\varphi(d)}\right|\ll_{B}\frac{\operatorname{\mathcal{X}}}{(\log\operatorname{\mathcal{X}})^{B-2}}+\sqrt{y}D\ll\frac{\operatorname{\mathcal{X}}}{(\log\operatorname{\mathcal{X}})^{B-2}}$
@@
We only prove (2.21) as the proof of (2.20) is essentially identical. The case $\gamma\leq 1$ follows immediately from Proposition 2.7 so we assume $\gamma>1$ throughout. We split into three cases for $y$.
@@
$B-2\gamma+1\geq B_{\gamma}:=(B-1-\gamma^{2})/2$
for $B>\gamma^{2}+1$.
@@
In this case, the Brun–Titchmarsh theorem (see e.g. [26, Theorem 2]) gives
@@
for all $d\leq D$. Using these bounds and Lemma 2.6,
@@
$\sum_{\begin{subarray}{c}d\leq D\\ (d,\overline{\operatorname{\mathcal{P}}})=1\end{subarray}}E(d)\ll_{B}\frac{\operatorname{\mathcal{X}}}{(\log\operatorname{\mathcal{X}})^{B-2}}.$ (2.25)
Next, an application of the Brun–Titchmarsh theorem and Lemma 2.6 yields
@@
As one might expect, a small improvement is possible if one uses Hölder’s inequality in place of the Cauchy–Schwarz inequality. For example, provided $B>\gamma+3$}, one can apply Hölder’s inequality in the proof above, writing
$\sum_{\begin{subarray}{c}d\leq D\\ (d,\overline{\operatorname{\mathcal{P}}})=1\end{subarray}}f(d)E(d)\leq\Bigg({\sum_{\begin{subarray}{c}d\leq D\\ (d,\overline{\operatorname{\mathcal{P}}})=1\end{subarray}}f(d)^{r}E(d)}\Bigg)^{1/r}\Bigg(\sum_{\begin{subarray}{c}d\leq D\\ (d,\overline{\operatorname{\mathcal{P}}})=1\end{subarray}}E(d)\Bigg)^{1/s}.$
@@
$(r,s)=\left(\frac{\log(B-3)}{\log\gamma},\frac{\log(B-3)}{\log(B-3)-\log\gamma}\right).$
@@
$\widetilde{B}_{\gamma}=B-2-\frac{2(B-2)\log\gamma}{\log(B-3)}.$
So, for instance, if $B=11$ and $\gamma=2$ then $\widetilde{B}_{\gamma}=4$ which is greater than $B_{\gamma}=3$}. Nevertheless, we use $B_{\gamma}$ in the statement of Theorems 1.6 and 1.7 for simplicity.